# MindShift: Leveraging Large Language Models for Mental-States-Based Problematic Smartphone Use Intervention

### v1

@@ 7.1.4. Acceptance Rate across Different Strategies
(a) Strategy Acceptance Rate Grouped by Mental State. Significant differences compared to no strategy (green bar) are indicated above the X coordinate axis, while differences among strategies are below.
@@
Figure 9. Strategy Acceptance Rate across Different Mental States and Activities
@@ 7.1.5. Strategy Acceptance Rate across Different Mental States and Activities
As we show in Figure 2, each mental state has a different strategy mapping. Therefore, we also sought to derive insights regarding which strategies are most effective for users under different mental states (Figure 8(a)) and activities (Figure 8(b)). Friedman test and post hoc Wilcoxon signed-rank tests are employed to investigate the influence of different strategies on acceptance rate.
@@
For activity levels: When not engaging in activity, Comforting is significantly more effective than MindShift-Simple ($\Delta_{\textit{}}$=5.6%, p ¡ 0.05). When engaging in activity, there are no significant differences in all the strategies compared to MindShift-Simple.
@@ 7.2.1. Overall Usage Behavior
We count the number of app opening attempts for restricted apps. Figure 9(a) presents the opening frequency under three intervention techniques. MindShift-Simple (63.6±9.2) and MindShift (65.3±9.5) have lower opening frequency than Baseline (74.3±9.7). Compared to Baseline, MindShift reduces by 12.1% usage duration while MindShift-Simple reduces by 14.4%. However, a Friedman test does not show significance.
We also measure restricted app usage duration which is another important factor for phone overuse. As can be seen from Figure 9(b), participants have the lowest app usage duration in MindShift (1.11 ±0.7 hours) compared to the Baseline (1.23±0.7 hours) and MindShift-Simple (1.20 ± 0.8 hours). Compared to Baseline, MindShift reduces by 9.8% usage duration while MindShift-Simple reduces by 2.4%. We conduct a Friedman test and find a significant difference among different techniques (p¡0.05). A post-hoc Wilcoxon test shows that MindShift has a trend of declining compared to Baseline, with marginal significance (p¡0.1).
@@ 7.3. Subjective Report
7.3.1. Decrease of SAS Score
Both MindShift and MindShift-Simple decrease SAS Score significantly and MindShift achieves best. Figure 10(a) shows the results of the SAS scores during the intervention stages. MindShift exhibits the lowest SAS scores (35.2±10.2), followed by MindShift-Simple in the second position (37.6±10.1), with the Baseline intervention ranking the third (44.5±8.7) and the initial ranking the last (47.5±6.7). We conduct a Friedman test and the results show a significant difference(p¡0.001). Post-hoc Wilcoxon tests show that both MindShift and MindShift-Simple are significantly lower than the Baseline and initial scores (all ps¡0.01). This suggests that the two versions of MindShift have the potential to fundamentally alter individuals’ mobile phone usage behavior.
7.3.2. Increase of Self-efficacy
MindShift and MindShift-Simple increase Self-efficacy Score significantly. Figure 10(b) shows the results of the self-efficacy scores during the intervention stages. MindShift (26.7±6.1) and MindShift-Simple (26.6±6.8) exhibit higher scores compared to the Baseline intervention (24.1±5.6) and the initial (24.1±5.6). We conduct a Friedman test, and the results show a significant difference (p¡0.001). The post-hoc Wilcoxon test shows that both MindShift and MindShift-Simple are significantly higher than the Baseline and initial scores (all ps¡0.05). This indicates that persuasive techniques have the potential to significantly enhance individuals’ self-efficacy, which can result in overcoming excessive mobile phone usage. In contrast, conventional reminders may not achieve this goal.
@@ 7.3.3. Subjective Comments
During the exit interview, participants generally had a positive experience. One participant said, “I can actually feel that my recent dependency on the phone has decreased” (P17). Another participant felt using the app shifted them to self-improvement tasks instead of mindlessly using the smartphone, “Now, when I have nothing to do, I tend to do other things, like learning vocabulary, instead of aimlessly browsing my phone” (P10). Additionally, participants had positive comments on our design which includes the mental states factor during the intervention, “I feel that its suggestions align well with my emotional state at that time.” (P25). They also valued the “Understanding” and “Comforting” strategies, saying “It tells me that I’m not the only one experiencing these painful emotions, which helps me feel better” (P22).
Despite the majority positive comments, a small number of participants expressed their dissatisfaction with MindShift. Some participants found the persuasive message to be “a bit stiff and templated” (P4, P6), and they believed that “they would develop tolerance as they repeatedly use” (P20). Some participants also mentioned privacy concerns. For example, P15 mentioned that MindShift-Simple’s ability to capture time and location made her uncomfortable.
@@ 7.4. Summary of Results
Overall, two versions of MindShift show significantly higher acceptance rates compared to the Baseline. MindShift has the highest acceptance (35.2%) and thumb-up rates (6.8%), and it significantly outperforms MindShift-Simple in both acceptance rates of generated persuasion content (8.1%) and on a per-round basis (6-14.7%). Furthermore, there exist strategies that significantly outperform MindShift-Simple (3.6-10.8%) in every mental state, suggesting that the strategies we design are meaningful. Regarding app usage behavior, two versions of MindShift lead to a decrease in overall app opening frequency (12.1-14.4% ) and usage duration (2.4-9.8%) and are significantly effective in reducing habitual use. MindShift achieves the best results in reducing SAS scores (34.7%) and increasing self-efficacy scores (10.7%). Statistical analysis indicates that two versions of MindShift reduce SAS scores and increase self-efficacy scores significantly while Baseline (traditional pop-up reflections) does not. This suggests that MindShift has the potential to profoundly transform human behavior with enduring effects. Finally, users’ subjective comments also confirm a perceived reduction in smartphone dependency and an inclination to continue using MindShift.
@@ 8. DISCUSSION
In this section, we discuss the MindShift’s novelty in contrast to previous intervention techniques and discuss the limitations and future work.
@@ 8.1. The Roles of Users’ Phone Use Purpose and Mental States in Smartphone Intervention
Most previous intervention techniques are designed to initiate interventions based on the amount of time and frequency of smartphone usage. However, quantifying smartphone use just by considering time oversimplifies and ignores the underlying causes. People use smartphones for work, study, and relaxation. MindShift initiates persuasion only when users agree the current usage belongs to habitual use. Our intervention enhances users’ self-awareness of their habitual phone use behavior.
Additionally, certain mental states are linked to habitual smartphone use as a form of self-distraction. Although smartphone use serves as a coping mechanism for emotion fluctuations, studies show that habitually using them for escapism fails to effectively mitigate emotions (Duvenage et al., 2020). Using smartphones for emotional regulation can lead to problematic smartphone use behavior, potentially leading to severe psychological issues such as depression (Coyne et al., 2021; Zsido et al., 2021). MindShift aims to intervene in habitual smartphone usage triggered by specific mental states. Our goal is to reduce users’ problematic smartphone use and help individuals transition from avoidance-oriented coping to approach-oriented coping (Compas et al., 2001). Future work can also explore automatic detection of users’ phone use intention and mental states, simplifying the intervention process and improving user experience. However, it remains an open question on how skipping the self-reflection may impact the effectiveness of such a persuasion technique.
@@ 8.2. Towards Adaptive Persuasion Intervention
MindShift generates dynamic and personalized persuasion content by combining information such as users’ physical contexts, mental states, and other behaviors. However, several participants still mentioned that the generated content sometimes could be “stiff and templated”. This may be attributed to the limited prompt templates. Although the content generated by the LLM varies, the main theme is guided by our prompts, which could limit the variation of persuasion content. To improve, we suggest to integrate user feedback into the system. Currently, MindShift supports a simple thumb-up and thumb-down feedback mechanism. Even with such simple information, we could establish a human-in-the-loop setup to fine-tune content, aligning better with user preferences. Moreover, future work can also consider collecting more comprehensive feedback from users. Users could customize the language style generated by an LLM, which can be coupled with adaptive algorithms such as reinforcement learning to achieve a more intelligent just-in-time adaptive intervention (JITAI) system that evolves with users  (Goldstein et al., 2017).
8.3. Ethical Concerns and Risk of Using LLM for Persuasion
Although MindShift performs well in reducing smartphone use frequency, there are important ethical concerns we want to highlight, especially about the application of LLMs for persuasion. Despite our intentionally devised prompts given to the LLM, it is not possible for us to guarantee that the generated persuasive messages are always safe. For example, if users access a blacklisted shopping app, LLM might include unwelcome commercial ideas in the message, causing user aversion and rejection. Meanwhile, because of LLM’s limited capability in understanding human psychology and society (Xu et al., 2023), it may evoke goals or recommend habits at the wrong time. For instance, if users are already stressed due to their life objectives (e.g., struggling with academic stress), additional reminders could exacerbate the stress or even cause harm. Recent research in the field of AI alignment aims to ensure that AI systems behave in ways that are safe, beneficial, and aligned with human values and objectives (Gabriel, 2020). Real-world deployment is needed to integrate these techniques into the intervention.

### v2

@@ 7.1.3. Thumb-up Rate of Interventions
(Top) A bar plot of the feedback proportion of two versions of MindShift. The X-axis is the two feedback categories. From the left to the right are Thumb-Up, and Thumb-Down, and each category has two bars, the left is MindShift-Simple while the right is MindShift. The Y-axis is the proportion rate. The thumb-up proportion for MindShift-Simple is 2.2% while MindShift is 6.8%. The thumb-down proportion for MindShift-Simple is 1.6% while MindShift is 2.3%. (Bottom) A bar plot of the persuasion acceptance rate of different strategies. The Y-axis is the acceptance rate. The X-axis is the strategy. From the left to the right are understanding, comforting, evoking, scaffolding habits, and MindShift-Simple, they have acceptance rates of 20.9%, 15.3%, 22.3%, 16.9%, and 12.0% respectively. Each strategy has a higher acceptance rate than MindShift-Simple.
@@ 7.1.4. Acceptance Rate across Different Strategies
Figure 9. Strategy Acceptance Rate Grouped by Different Mental States
(a) Strategy Acceptance Rate Grouped by Mental State. Significant differences compared to no strategy (green bar) are highlighted in red, while differences among strategies are indicated in black.
@@
(Top) A bar plot of the strategy acceptance rate grouped by 3 mental states. The Y-axis is the acceptance rate. The X-axis is the mental state. From the left to the right are stress, inertia, and boredom. Each mental state has five bars, from left to right are understanding, comforting, evoking, scaffolding habits, and MindShift-Simple. Compared with MindShift-Simple, under the mental state of ”Stress”, every strategy has a higher acceptance rate. Under the mental state of Inertia” and ”Boredom”, Evoking is higher. (Bottom) A bar plot of the strategy acceptance rate grouped by 2 activity states. The Y-axis is the acceptance rate. The X-axis is the activity state. From the left to the right are not engaging in activity and engaging in activity. Each activity state has five bars, from left to right are understanding, comforting, evoking, scaffolding habits, and MindShift-Simple. Compared with MindShift-Simple, when not engaging in activity, understanding and comforting have a higher acceptance rate. when engaging in activity, all strategies have a higher acceptance rate.
Figure 9. Strategy Acceptance Rate Grouped by Different Mental States
@@ 7.1.5. Strategy Acceptance Rate across Different Mental States and Activities
As we show in Figure 3, each mental state has a different strategy mapping. Therefore, we also seek to derive insights regarding which strategies are most effective for users under different mental states (Figure 8(a)) and activities (Figure 8(b)). Friedman test and post hoc Wilcoxon signed-rank tests are employed to investigate the influence of different strategies on acceptance rate.
@@
For activity levels: when not engaging in activity, Comforting is significantly more effective than MindShift-Simple ($\Delta_{\textit{}}$=5.6%, p ¡ 0.05); when engaging in activity, there are no significant differences in all the strategies compared to MindShift-Simple.
@@ 7.2.1. Overall Usage Behavior
We count the number of app opening attempts for restricted apps. Figure 9(a) presents the opening frequency (daily open count) under three intervention techniques. MindShift-Simple (63.6±9.2) and MindShift (65.3±9.5) have lower opening frequency than Baseline (74.3±9.7). Compared to Baseline, MindShift reduces by 12.1% usage duration while MindShift-Simple reduces by 14.4%. However, a Friedman test does not show significance.
We also measure restricted app usage duration, another important factor for phone overuse. As can be seen from Figure 9(b), participants have the lowest app usage duration in MindShift (1.11 ±0.7 hours) compared to the Baseline (1.23±0.7 hours) and MindShift-Simple (1.20 ± 0.8 hours). Compared to Baseline, MindShift reduces by 9.8% usage duration while MindShift-Simple reduces by 2.4%. We conduct a Friedman test and find a significant difference among different techniques (p¡0.05). A post-hoc Wilcoxon test shows that MindShift has a trend of declining compared to Baseline, with marginal significance (p¡0.1).
To validate the assumption that the intent and mental state remain stable in one unlock session across different apps in Sec. 4.3, we analyze the duration of users’ intent and mental states. The results show that the median duration of a mental state is 5 hours (third quartile 14.5 hours). The median duration for an intent (changing from habitual use to other intents) is 37 minutes (third quartile 60 minutes). This validates our hypothesis that intent and mental state are stable during one habitual usage session.
Figure 10. App Usage Behavior
@@
(Top) A bar plot of the app opening frequency of all restricted apps of the three techniques. The X-axis is the three techniques. From the left to the right are Baseline, MindShift-Simple, and MindShift. The Y-axis is the opening frequency per day. MindShift-Simple (63±9.2) and MindShift (65±9.5) had lower opening frequency than Baseline (74±9.7). (Bottom) A bar plot of the app usage duration of all restricted apps of the three techniques. The X-axis is the three techniques. From the left to the right are Baseline, MindShift-Simple, and MindShift. The Y-axis is the usage duration per day. MindShift (1.11 ±0.7 hours) has the lowest app usage duration compared to the Baseline (1.23±0.7 hours) and MindShift-Simple(1.20 ± 0.8 hours).
@@ 7.2.2. Habitual Usage Behavior
To confirm the suitability of the 2-minute intervention interval, we analyze data on users’ habitual usage duration during the Baseline phase (unaffected by subsequent interventions). Analysis shows that 75% of users spent about 4 minutes in a single habitual use, supporting our design choice of the 2-min intervention interval.
@@ 7.3. Subjective Report
Figure 11. Subjective Scales Report
@@
(Top) A bar plot of the SAS score of the four intervention stages. The X-axis is the four stages. From the left to the right are Before Intervention, Baseline, MindShift-Simple, and MindShift. The Y-axis is the SAS Score. MindShift exhibited the lowest SAS scores (35.2±10.2), followed by MindShift-Simple in the second position (37.6±10.1), with the Baseline intervention ranking the third (44.5±8.7) and the initial ranking the last (47.5±6.7). (Bottom) A bar plot of the Self-efficacy score of the four intervention stages. The X-axis is the four stages. From the left to the right are Before Intervention, Baseline, MindShift-Simple, and MindShift. The Y-axis is the Self-efficacy Score. MindShift(26.7±6.1) and MindShift-Simple(26.6±6.8) exhibited higher scores compared to the Baseline intervention (24.1±5.6) and the initial (24.1±5.6).
@@
7.3.1. Decrease of SAS Score
MindShift and MindShift-Simple decrease SAS score significantly and MindShift achieves best. Figure 10(a) shows the results of the SAS scores during the intervention stages. MindShift exhibits the lowest SAS scores (35.2±10.2), followed by MindShift-Simple in the second position (37.6±10.1), with the Baseline intervention ranking the third (44.5±8.7) and the initial ranking the last (47.5±6.7). This indicates that MindShift and MindShift-Simple reduce SAS scores by 34.7 and 25.8%. The results of a Friedman test show a significant difference (p¡0.001). Post-hoc Wilcoxon tests show that both MindShift and MindShift-Simple are significantly lower than the Baseline and initial scores (all ps¡0.01). This suggests that the two versions of MindShift have the potential to fundamentally alter individuals’ mobile phone usage behavior.
7.3.2. Increase of Self-efficacy Score
MindShift and MindShift-Simple increase Self-efficacy score significantly. Figure 10(b) shows the results of the self-efficacy scores during the intervention stages. MindShift (26.7±6.1) and MindShift-Simple (26.6±6.8) exhibit higher scores compared to the Baseline intervention (24.1±5.6) and the initial (24.1±5.6). This indicates that MindShift and MindShift-Simple increase self-efficacy scores by 10.7% and 10.4%. Friedman test shows a significant difference (p¡0.001). The post-hoc Wilcoxon test shows that both MindShift and MindShift-Simple are significantly higher than the Baseline and initial scores (all ps¡0.05). This indicates that persuasive techniques have the potential to enhance individuals’ self-efficacy, which can result in overcoming excessive mobile phone usage. In contrast, conventional reminders may not achieve this goal.
@@ 7.3.3. Subjective Comments
During the exit interview, participants generally had a positive experience. One participant said, “I can feel that my recent dependency on the phone has decreased” (P17). One participant felt using the app shifted them to self-improvement tasks instead of mindlessly usage, “Now, when I have nothing to do, I tend to do other things, like learning vocabulary, instead of aimlessly browsing my phone” (P10). Another participant was willing to use it longer, “I’m a little sad with the disappearance of pop-ups after the experiment. If possible, I would like to keep using it” (P3). Additionally, participants had positive comments on MindShift which includes the mental states factor, “I feel that its suggestions align well with my emotional state at that time.” (P25). They also valued the “Understanding” and “Comforting” strategies, saying “It tells me that I’m not the only one experiencing these painful emotions, which helps me feel better” (P22).
Despite the majority of positive comments, a small number of participants expressed their dissatisfaction with MindShift. Some participants found the persuasive message to be “a bit stiff and templated” (P4, P6), and they believed that “they would develop tolerance as they repeatedly use” (P20). Some participants also mentioned privacy concerns. P15 mentioned that MindShift-Simple’s ability to capture time and location made her uncomfortable. Moreover, participants’ preference for linguistic characteristics is highly personal. Some felt harsh ones were more useful, “Gentle tone doesn’t work for me, I wish the words could be harsher” (P8). Some preferred data proof than pure textual reasoning, “It’s intuitive to tell me how long I have used my phone today directly. The number is very eye-catching” (P9). This suggests the future direction of personalized persuasive content design. We have more discussion in Sec. 8.2.
@@ 7.4. Summary of Results
Overall, two versions of MindShift show significantly higher acceptance rates compared to the Baseline. MindShift has the highest acceptance (45.1% for session-based and 35.2% for pop-up-based) and thumb-up rates (6.8%), and it statistically significantly outperforms MindShift-Simple in both acceptance rates of generated persuasion content (8.1%) and on a per-round basis (6-14.7%). Furthermore, there exist strategies that significantly outperform MindShift-Simple (3.6-10.8%) in every mental state, suggesting that the strategies we design are meaningful. Furthermore, MindShift and MindShift-Simple lead to a decrease in overall app opening frequency (12.1-14.4%) and usage duration (9.8-2.4%) and are significantly effective in reducing habitual use. MindShift and MindShift-Simple also reduce SAS scores by 34.7-25.8% and increase self-efficacy scores by 10.7-10.4% statistically significantly while Baseline does not. This suggests that MindShift has the potential to profoundly transform human behavior with enduring effects. Finally, users’ subjective comments also confirm a perceived reduction in smartphone dependency and an inclination to continue using MindShift.
@@ 8. DISCUSSION
In this section, we discuss MindShift’s novelty in contrast to previous intervention techniques (Sec 8.1), future work (Sec 8.2), the potential of leveraging LLMs for behavior change (Sec 8.3), and the limitations (Sec 8.4).
@@ 8.1. The Roles of Users’ Phone Use Purpose and Mental States in Smartphone Intervention
Most previous intervention techniques initiate interventions based on the amount of time and frequency of smartphone usage. However, quantifying smartphone use just by time oversimplifies and ignores the underlying causes. People use smartphones for work, study, and relaxation, as long as for meaningful reasons, the time is not a true problem. As Lukoff et al. found, even if participants didn’t reduce their screen time, the intervention could make them feel better in the sense of agency and goal alignment, indicating users prioritize the quality of time over quantification  (Lukoff et al., 2023). MindShift initiates persuasion only when users recognize their current usage as habitual, aiming to enhance users’ self-awareness of their habitual phone use behavior.
Additionally, certain mental states are linked to habitual smartphone use as a form of self-distraction. Although smartphone use serves as a coping mechanism for emotion fluctuations, studies show that habitually using them for escapism fails to effectively mitigate emotions (Duvenage et al., 2020). Using smartphones for emotional regulation can lead to problematic smartphone use behavior, potentially leading to severe psychological issues such as depression (Coyne et al., 2021; Zsido et al., 2021). MindShift aims to intervene in habitual smartphone usage triggered by specific mental states. Our goal is to reduce users’ problematic smartphone use and help individuals transition from avoidance-oriented coping to approach-oriented coping (Compas et al., 2001).
@@ 8.2. Towards Adaptive Persuasion Intervention
MindShift generates dynamic and personalized persuasion content by combining information such as users’ simple physical contexts, mental states, and other behaviors. However, several participants still mentioned that the LLM-generated content sometimes could be “stiff and templated”. This may be attributed to the limited prompt templates. Although the content generated by the LLM varies, the main theme is guided by our prompts, which could limit the variation of persuasion content. To improve, we suggest integrating user feedback into the system for more adaptive intervention. Currently, MindShift supports a simple thumb-up and thumb-down feedback mechanism. Even with such simple information, we could establish a human-in-the-loop setup to fine-tune content, aligning better with user preferences. Another aspect is to include more diverse behavior features captured by passive sensors on smartphones and wearables (Xu et al., 2022a; Meegahapola et al., 2023; He et al., 2020).
Moreover, future work can also consider collecting more comprehensive feedback from users. Users could customize the language style generated by an LLM, which can be coupled with adaptive algorithms such as reinforcement learning to achieve a more intelligent just-in-time adaptive intervention (JITAI) system that evolves with users  (Goldstein et al., 2017; Orzikulova et al., 2024).
8.3. Leveraging LLMs for Behavior Change
8.3.1. Advantages of Using LLMs for Behavior Change
Our study sheds light on the possibility of leveraging LLMs to change user behavior by influencing human cognition. Previous efforts in this field often hinge on users’ ability to self-reflect and self-persuade, limited by the narrow scope of sentence databases used (Xu et al., 2022b). LLMs break this barrier, offering a broader range of persuasive strategies. They can generate adaptive and diverse persuasive content, tailored to the individual’s context. In our study, we observed notable changes in cognition: participants’ smartphone addiction scale scores dropped, and self-efficacy scores rose after the study (Figure 10(a)). This suggests a potential for long-term behavioral change, which we aim to explore further in our future work.
We also want to highlight that MindShift is just one example of the possibilities in this domain. Future research could integrate LLMs for more dynamic interventions, such as self-affirmation content generation in the typing intervention (Xu et al., 2022b) and personalized visualizations (Hiniker et al., 2016a). Moreover, our methodology can potentially be expanded to other domains, such as smoking or alcohol cessation, eating diet, and physical activity promotion, where LLMs can be used to generate context-aware dynamic persuasive content for a specific well-being goal. We envision our work can inspire a number of creative LLM-powered intervention techniques in the future.
8.3.2. Design Implication for Using LLMs in Other Behavior Change Domains
Based on our findings in the study, we extract three design implications of using LLMs for behavior change in various domains.
First, investigating why people behave in certain ways is the basis of any intervention design, especially when LLMs can utilize such insights when generating persuasion. Our study explores the psychological factors behind habitual smartphone use. MindShift leveraging those factors outperformed MindShift-Simple only considering physical factors (see in Sec. 7.1.1 to 7.1.3).
Second, context is crucial for LLMs to generate dynamic, tailored persuasive messages. Our examples in Table 2 showcase the importance of user contexts for content generation. In our study, some user contexts, like mental states, cannot be detected automatically but depend on users’ self-reports, which can be improved in future design. Past work has explored how to use smartphone usage data and machine learning to predict boredom and stress (Pielot et al., 2015; Lekkas et al., 2022; Ciman and Wac, 2016; Stütz et al., 2015; Meegahapola et al., 2023), and there have been researches using physiological measuring instruments to learn mental states from biosignals (Seo et al., 2019b, a; Xu et al., 2019). With the development of more smart and wearable technologies, there is the potential to track users’ mental states automatically (Hickey et al., 2021; Gedam and Paul, 2021). This can simplify the intervention process and improve user experience. However, it remains an open question on how skipping self-reflection may impact the effectiveness of such a persuasion technique.
Last, crafting a suitable prompt is crucial for effectively incorporating expert knowledge into LLM generation. This often involves multiple attempts and adjustments to ensure the generated content aligns with expectations. While not the main contribution of our work, we conducted extensive iterations to ensure the appropriateness of the persuasive content. We have more discussion on the ethical concerns if prompt engineering is not done properly in the next paragraph. We refer future developers to recent studies, such as EmotionPrompt (Li et al., 2023), for more comprehensive guidance on enhancing LLM outputs.
8.3.3. Ethical Concerns and Risk of Using LLMs for Behavior Change
Although MindShift performs well in changing problematic smartphone use, there are important ethical concerns we want to highlight about the risk of using LLMs for behavior change.
Despite carefully crafted prompts, developers face challenges ensuring the constant safety of generated persuasive messages. For example, while we fixed hallucination for our experiment, it is one of the biggest concerns in LLMs and can still possibly occur in real-world deployment (Ji et al., 2023; Ferrer et al., 2021). Additionally, although we didn’t encounter it in our experiment, LLMs can generate dangerous content, such as abusive and discriminatory sentences. Furthermore, there is still room to improve LLMs’ understanding of the nuances of human mental states (Xu et al., 2023). For instance, if users of MindShift are already stressed due to their life objectives (e.g., struggling with academic stress), additional reminders of these goals could exacerbate the stress or even cause harm. More future work is needed to improve safety and reduce the risk before we deploy LLMs for large-scale intervention studies.